Joe is 10 yrs. older than James. In 8 yrs. twice Joe,s age will equal three times James age. How old is each now?
Answer by ankor@dixie-net.com(22740)

Joe is 10 yrs. older than James.
Jo = Ja + 10
:
In 8 yrs. twice Joe,s age will equal three times James age.
2(Jo+8) = 3(Ja+8)
2Jo + 16 = 3Ja + 24
2Jo = 3Ja + 24 - 16
2Jo = 3Ja + 8
:
How old is each now?
From the 1st equation, replace Jo with (Ja+10)
2(Ja+10) = 3Ja + 8
2Ja + 20 = 3Ja + 8
20 - 8 = 3Ja - 2Ja
12 = Ja
:
Ja is 12 yrs old, therefore
Jo = 12 + 10
Jo = 22 yrs
:
:
Check solutions in the statement:
"In 8 yrs. twice Joe,s age will equal three times James age."
2(22 + 8) = 3(12+8)
2(30) = 3(20)